OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 2x + y = 5 x + 3y = 0

OpenStudy (callisto):

2x + y = 5 -(1) x + 3y = 0 => x=-3y -(2) Sub. (2) into (1) 2 (-3y) +y =5 y=... Sub value of y into (2), you can get x.

OpenStudy (anonymous):

y=10

OpenStudy (anonymous):

2x + y= 5 - (eqn.1) x+3y = 0 - (eqn.2) Taking eqn.2 and multiply it by 2, x+3y=0 2x+6y=0 - (eqn.3) Take, eqn.3 and minus eqn.1 2x+6y- (2x+y) = 0-5 5y = -5 y=-1 Sub y=-1 to eqn.1, 2x-1 = 5 2x=6 x=3

OpenStudy (callisto):

Not really 2(-3y) +y =5 -6y +y =5 -5y = 5 y = ?

OpenStudy (anonymous):

oh i get it..yeah its -1